Exercise 11-21 Cash dividends, treasury stock, and statement of retained earnings LO C3, P2, P3 Alexander Corporation reportsJournal entry worksheet 2 3 4 5 6 7 8 Record the purchase of 3,300 shares of its own common stock for $25 cash per share. Notindicated by a minus sign.) ALEXANDER CORPORATION Statement of Retained Earnings For Year Ended December 31ALEXANDER CORPORATION Stockholders Equity Section of the Balance Sheet December 31 Total stockholders equity
Answer #1
1
Date General Journal Debit Credit
2-Jan Treasury stock, common 82,500 =3300*25
Cash 82,500
7-Jan Retained earnings 44,550 =(33000-3300)*1.50
Common dividend payable 44,550
28-Feb Common dividend payable 44,550
Cash 44,550
9-Jul Cash 39,600 =1320*30
Treasury stock, common 33,000 =1320*25
Paid-in capital, treasury stock 6,600
27-Aug Cash 33,000 =1650*20
Paid-in capital, treasury stock 6,600
Retained earnings 1,650
Treasury stock, common 41,250 =1650*25
9-Sep Retained earnings 65,340 =(33000-3300+1320+1650)*2
Common dividend payable 65,340
22-Oct Common dividend payable 65,340
Cash 65,340
31-Dec Income summary 55,000
Retained earnings 55,000
2
ALEXANDER CORPORATION
Statement of Retained Earnings
For Year Ended December 31
Retained earnings, December 31, 2018 350,000
Add: Net income 55,000
405,000
Less: Cash dividends declared (109,890)
Less: Treasury stock reissuances (1,650)
Retained earnings, December 31, 2019 293460
3
ALEXANDER CORPORATION
Stockholders’ Equity Section of the Balance Sheet
December 31
Common stock - $25 par value 825,000
Paid-in capital in excess of par value, common stock 66,000
Retained earnings 293,460
Less: Cost of treasury stock (8,250)
Total stockholders’ equity 1,176,210